OpenStudy (anonymous):

In a school of 200 students, 80 students are in the band, 160 students are on sports teams, and 60 students participate in both activities. How many students in the school are neither in band nor on a sports team?

OpenStudy (anonymous):

\[200-(80+160+60)=...\] that's how you get the answer i think... im only going to 6th grade on august 1st...

OpenStudy (anonymous):

let A be the set of students in the band, B be the set of students in sports team. then A=80, B=160. A intersection B= 60. no. of the students that are neither in band nor in sports team is given by the complement of a union B. now, a union B = A + B -A intersection B = 80 + 160 - 60 = 180. so the answer is 200 - 180 =20.

OpenStudy (anonymous):

oooooo okay Ananya